OpenStudy (kendricklamar2014):

Find the simplified form of the expression. Give your answer in Scientific Notation. (5 * 10^6)(3*10^5) A. 1.5 * 10^11 B. 1.5 * 10^12 C. 15 * 10^11 D. 1.5 * 10^10

OpenStudy (kendricklamar2014):

@pooja195 @SolomonZelman

OpenStudy (michele_laino):

please note that you have to multiply this: \[5 \cdot 3 \cdot {10^6} \cdot {10^5}\]

OpenStudy (michele_laino):

what is 5*3=...?

OpenStudy (kendricklamar2014):

5*3 = 15

OpenStudy (michele_laino):

ok!

OpenStudy (michele_laino):

and what is: \[{10^6} \cdot {10^5} = ...?\]

OpenStudy (kendricklamar2014):

10^6 * 10^5 = 10^11

OpenStudy (michele_laino):

ok! so yur result is: \[15 \cdot {10^{11}}\]

OpenStudy (kendricklamar2014):

So the answer would be C

OpenStudy (michele_laino):

please now, keep in mind that \[15 = 1.5 \cdot {10^1}\]

OpenStudy (kendricklamar2014):

So, would the answer still be C or no...??

OpenStudy (michele_laino):

no, I don't think, since if I make this substitution: \[15 = 1.5 \cdot {10^1}\] I get: \[5 \cdot 3 \cdot {10^6} \cdot {10^5} = 1.5 \cdot {10^1} \cdot {10^{11}}\]

OpenStudy (kendricklamar2014):

So it would come out to be 1.5 * 10^12

OpenStudy (michele_laino):

that's right!

OpenStudy (kendricklamar2014):

Thanks, can you help me with one more??

OpenStudy (michele_laino):

ok!

OpenStudy (anonymous):

Really? I thought it would be D.

OpenStudy (michele_laino):

please since 15 = 1.5 * 10^1, then 15*10^11= 1.5 * 10 * 10^11 = 1.5 *10^12 @taylorswift131313

OpenStudy (anonymous):

Ok, can you give me a medal @Michele_Laino? I gave you one.

OpenStudy (anonymous):

thank you anyway